# Lotus: Diffusion-based Visual Foundation Model for High-quality Dense Prediction

### v2

@@ 2.3 Monocular Depth and Normal Prediction
Monocular depth and normal prediction are two crucial dense prediction tasks. Solving them typically demands comprehensive scene understanding capability. Starting from Eigen et al. (2014), early CNN-based methods for depth prediction, such as Fu et al. (2018), Lee et al. (2019), Yuan et al. (2022), focus only on specific domains. Subsequently, in pursuit of a generalizable depth estimator, many methods expand model capacity and train on larger and more diverse datasets, such as DiverseDepth (Yin et al., 2021a) and MiDaS (Ranftl et al., 2020). DPT (Ranftl et al., 2021) and Omnidata (Eftekhar et al., 2021) are further proposed based on vision transformer (Ranftl et al., 2021), significantly enhancing performance. LeRes (Yin et al., 2021b) and HDN (Zhang et al., 2022) further introduce novel training strategies and multi-scale depth normalization to improve predictions in detailed areas. More recently, the DepthAnything series (Yang et al., 2024a; b) and Metric3D series (Yin et al., 2023; Hu et al., 2024) collect and leverage millions of training data to develop more powerful estimators. Normal prediction follows the same trend. Starting with the early CNN-based methods like OASIS (Chen et al., 2020), EESNU (Bae & Davison, 2021) and Omnidata series (Eftekhar et al., 2021; Kar et al., 2022) expand the model capacity and scale up the training data. Recently, DSINE (Bae & Davison, 2024) achieves SoTA performance by rethinking inductive biases for surface normal estimation. In our paper, we focus on leveraging pre-trained diffusion priors to enhance zero-shot dense predictions, rather than expanding model capacity or relying on large training data, which avoids the need for intensive resources and computation.
@@ 3 Preliminaries
Diffusion Formulation for Dense Prediction. Following Ke et al. (2024) and Fu et al. (2024), we also formulate dense prediction as an image-conditioned annotation generation task based on Stable Diffusion (Rombach et al., 2022), which performs the diffusion process in low-dimensional latent space for computational efficiency. First, there is a pair of auto-encoders $\{\mathcal{E}(\cdot),\mathcal{D}(\cdot)\}$ trained to map between RGB space and latent space, i.e., $\mathcal{E}(\textbf{x})=\mathbf{z^{x}}$, $\mathcal{D}(\mathbf{z^{x}})\approx\textbf{x}$. The auto-encoder also maps between dense annotations and latent space effectively, i.e., $\mathcal{E}(\textbf{y})=\mathbf{z^{y}}$, $\mathcal{D}(\mathbf{z^{y}})\approx\textbf{y}$ (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024). Following Ho et al. (2020), Stable Diffusion establishes a pair of forward nosing and reversal denoising processes in latent space. In forward process, Gaussian noise is gradually added at levels $t\in\left[1,T\right]$ into sample $\mathbf{z^{y}}$ to obtain the noisy sample $\mathbf{z^{y}_{t}}$:
@@ 4 Methodology
We start our analysis by directly adapting the original diffusion formulation with minimal modifications as illustrated in Fig. 4. We call this starting point as “Direct Adaptation [^1] Details of Direct Adaptation will be provided in Sec. B of supplementary materials.. Direct Adaptation is optimized using the standard diffusion objective as formulated in Eq. 2 (first row) and inferred by standard multi-step DDIM sampler. As shown in Tab. 3, Direct Adaptation fails to achieve satisfactory performance. In following sections, we will systematically analyze the key factors that affect adaptation performance step by step: parameterization types (Sec. 4.1); number of time-steps (Sec. 4.2); and the novel detail preserver (Sec. 4.3).
@@ 4.1 Parameterization Types
The type of parameterization is a vital configuration, it not only determines the loss function discussed in Sec. 3, but also influences the inference process (Eq. 3). During inference, the predicted clean sample $\mathbf{\hat{z}^{y}_{\tau}}$, a key component in Eq. 3, is calculated according to different parameterizations  [^2] The latest parameterization, $v$-prediction, combines $\epsilon$-prediction and $x_{0}$-prediction, producing results that are intermediate between the two. Please see the Sec. D of supplementary materials for more details..
@@ 5.1 Experimental Settings
Implementation details. We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), with text conditioning disabled. During training, we fix the time-step $t=1000$. For more details, please see the Sec. A.1 of supplementary materials.
@@
Evaluation Datasets and Metrics. ① For zero-shot affine-invariant depth estimation, we evaluate Lotus on NYUv2 (Silberman et al., 2012), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and ETH3D (Schops et al., 2017) using absolute mean relative error (AbsRel), and also report $\delta 1$ and $\delta 2$ values. ② For surface normal prediction, we employ NYUv2, ScanNet, iBims-1 (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets, reporting mean angular error (m.) as well as the percentage of pixels with an angular error below $11.25^{\circ}$ and $30^{\circ}$. Please see the Sec. A.2 of supplementary materials for further details on the evaluation datasets and metrics.
@@ 5.2 Quantitative Comparisons
① For depth estimation (Tab. 1), Lotus-G achieves the best performance compared to all generative baselines. Notice that we only require single step denoising process, also significantly boosting the inference speed as shown in Fig. 3. Lotus-D also performs well, only slightly inferior to DepthAnything. However, it is worthy to notice that Lotus is trained on only 0.059M images compared to DepthAnything’s 62.6M images. ② For normal estimation (Tab. 2), both Lotus-G and Lotus-D outperform all other generative and discriminative methods on zero-shot surface normal estimation with significant margins. Please see the Sec. E supplementary materials for Qualitative Comparisons.
@@
-- Noise Input (Lotus-D) 59K 5.311 96.733 99.186 9.662 91.637 98.643 6.757 95.382 98.992 5.786 96.339 99.136

### v4

@@ 2.3 Monocular Depth and Normal Prediction
Monocular depth and normal prediction are two crucial dense prediction tasks. Solving them typically demands comprehensive scene understanding capability. Starting from (Eigen et al., 2014), early CNN-based methods for depth prediction, such as (Fu et al., 2018; Lee et al., 2019; Yuan et al., 2022), focus only on specific domains. Subsequently, in pursuit of a generalizable depth estimator, many methods expand model capacity and train on larger and more diverse datasets, such as DiverseDepth (Yin et al., 2021a) and MiDaS (Ranftl et al., 2020). DPT (Ranftl et al., 2021) and Omnidata (Eftekhar et al., 2021) are further proposed based on vision transformer (Ranftl et al., 2021), significantly enhancing performance. LeRes (Yin et al., 2021b) and HDN (Zhang et al., 2022) further introduce novel training strategies and multi-scale depth normalization to improve predictions in detailed areas. More recently, the DepthAnything series (Yang et al., 2024a; b) and Metric3D series (Yin et al., 2023; Hu et al., 2024) collect and leverage millions of training data to develop more powerful estimators. Normal prediction follows the same trend. Starting with the early CNN-based methods like OASIS (Chen et al., 2020), EESNU (Bae & Davison, 2021) and Omnidata series (Eftekhar et al., 2021; Kar et al., 2022) expand the model capacity and scale up the training data. Recently, DSINE (Bae & Davison, 2024) achieves SoTA performance by rethinking inductive biases for surface normal estimation. In our paper, we focus on leveraging pre-trained diffusion priors to enhance zero-shot dense predictions, rather than expanding model capacity or relying on large training data, which avoids the need for intensive resources and computation.
@@ 3 Preliminaries
Diffusion Formulation for Dense Prediction. Following (Ke et al., 2024) and (Fu et al., 2024), we also formulate dense prediction as an image-conditioned annotation generation task based on Stable Diffusion (Rombach et al., 2022), which performs the diffusion process in low-dimensional latent space for computational efficiency. First, there is a pair of auto-encoders $\{\mathcal{E}(\cdot),\mathcal{D}(\cdot)\}$ trained to map between RGB space and latent space, i.e., $\mathcal{E}(\textbf{x})=\mathbf{z^{x}}$, $\mathcal{D}(\mathbf{z^{x}})\approx\textbf{x}$. The auto-encoder also maps between dense annotations and latent space effectively, i.e., $\mathcal{E}(\textbf{y})=\mathbf{z^{y}}$, $\mathcal{D}(\mathbf{z^{y}})\approx\textbf{y}$ (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024). Following (Ho et al., 2020), Stable Diffusion establishes a pair of forward nosing and reversal denoising processes in latent space. In forward process, Gaussian noise is gradually added at levels $t\in\left[1,T\right]$ into sample $\mathbf{z^{y}}$ to obtain the noisy sample $\mathbf{z^{y}_{t}}$:
@@ 4 Methodology
We start our analysis by directly adapting the original diffusion formulation with minimal modifications as illustrated in Fig. 4. We call this starting point as “Direct Adaptation [^1] Details of Direct Adaptation will be provided in Sec. B of the supplementary materials.. Direct Adaptation is optimized using the standard diffusion objective as formulated in Eq. 2 (first row) and inferred by standard multi-step DDIM sampler. As shown in Tab. 3, Direct Adaptation fails to achieve satisfactory performance. In following sections, we will systematically analyze the key factors that affect adaptation performance step by step: parameterization types (Sec. 4.1); number of time-steps (Sec. 4.2); and the novel detail preserver (Sec. 4.3).
@@ 4.1 Parameterization Types
The type of parameterization is a vital configuration, it not only determines the loss function discussed in Sec. 3, but also influences the inference process (Eq. 3). During inference, the predicted clean sample $\mathbf{\hat{z}^{y}_{\tau}}$, a key component in Eq. 3, is calculated according to different parameterizations  [^2] The latest parameterization, $v$-prediction, combines $\epsilon$-prediction and $x_{0}$-prediction, producing results that are intermediate between the two. Please see Sec. D of the supplementary materials for more details..
@@ 5.1 Experimental Settings
Implementation details. We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), with text conditioning disabled. During training, we fix the time-step $t=1000$. For more details, please see Sec. A.1 of the supplementary materials.
@@
Evaluation Datasets and Metrics. ① For zero-shot affine-invariant depth estimation, we evaluate Lotus on NYUv2 (Silberman et al., 2012), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and ETH3D (Schops et al., 2017) using absolute mean relative error (AbsRel), and also report $\delta 1$ and $\delta 2$ values. ② For surface normal prediction, we employ NYUv2, ScanNet, iBims-1 (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets, reporting mean angular error (m.) as well as the percentage of pixels with an angular error below $11.25^{\circ}$ and $30^{\circ}$. Please see Sec. A.2 of the supplementary materials for further details on the evaluation datasets and metrics.
@@ 5.2 Quantitative Comparisons
① For depth estimation (Tab. 1), Lotus-G achieves the best performance compared to all generative baselines. Notice that we only require single step denoising process, also significantly boosting the inference speed as shown in Fig. 3. Lotus-D also performs well, only slightly inferior to DepthAnything. However, it is worthy to notice that Lotus is trained on only 0.059M images compared to DepthAnything’s 62.6M images. ② For normal estimation (Tab. 2), both Lotus-G and Lotus-D outperform all other generative and discriminative methods on zero-shot surface normal estimation with significant margins. Please see Sec. E of the supplementary materials for Qualitative Comparisons.
@@
-- Noise Input (Lotus-D) 59K 5.334 96.729 99.198 9.334 92.813 98.795 6.846 95.290 98.899 5.982 96.287 99.087